# Anomalous dimensions from conformal field theory: generalized $\phi^{2n+1}$ theories

### v1

@@ Appendix F An OPE coefficient in generalized $\phi^{2n+1}$ theory ($n>1$)
Appendix G Generalized self-avoiding walks and loop-erased random walks
@@
As prominent models of random walks, we consider the self-avoiding walks and loop-erased random walks that play important roles in physics. Interestingly, they can be associated with the O($N$) symmetric $\phi^{4}$ theory. In the early 1970s, it was discovered that self-avoiding walks correspond to the $N\rightarrow 0$ limit deGennes:1972zz . Later, it was shown that loop-erased random walks are related to the $N\rightarrow-2$ limit duplantier1992loop ; Wiese:2018dow ; helmuth2020loop . In this appendix, we discuss these limits using the results from our previous paper Guo:2023qtt .
@@
In Guo:2023qtt , we studied the generalized $\phi^{2n}$ theory with O($N$) global symmetry.  [^32] It may be interesting to consider the $n>2$ counterpart of the special $N$ limits. The anomalous dimensions of $\phi$ at $n=2$ reads
@@
Moreover, the anomalous dimension of the symmetric traceless bilinear operator $\phi_{\text{T}}^{2}$ is
$\gamma_{\phi_{\text{T}}^{2}}=\frac{2}{N+8}\epsilon+O(\epsilon^{2})\,,$ (164)
@@
$D_{\text{F}}=d-\Delta_{\phi_{\text{T}}^{2}}=2-\frac{2}{N+8}\epsilon+O(\epsilon^{2})\,,$ (165)
where spacetime dimension is $d=4k-\epsilon$ and the fractal dimension $D_{\text{F}}$ is $k$-independent to order $\epsilon$. In particular, the limits $N\rightarrow 0$ and $N\rightarrow-2$ yield

### v2

@@ Appendix G The $N\rightarrow 0$ limit, logarithmic CFTs and logarithmic recombination
Appendix H Generalized self-avoiding walks, loop-erased random walks and their $\phi^{2n}$ generalizations
@@
As prominent models of random walks, we consider the self-avoiding walks and loop-erased random walks that play important roles in physics. Interestingly, they can be associated with the O($N$) symmetric $\phi^{4}$ theory. The self-avoiding walks correspond to the $N\rightarrow 0$ limit deGennes:1972zz  [^38] See Liu:2012ca for some Monte Carlo results for the loop model at various $N$ in three dimensions, including the case of $N=0$. and the loop-erased random walks are related to the $N\rightarrow-2$ limit Nienhuis:1982fx ; duplantier1992loop ; Wiese:2018dow ; Wiese:2019xmu ; helmuth2020loop . In this appendix, we discuss these limits and their $\phi^{2n}$ generalizations using the results from our previous paper Guo:2023qtt .
@@
In Guo:2023qtt , we studied the generalized $\phi^{2n}$ theory with O($N$) global symmetry. The anomalous dimensions of $\phi$ at $n=2$ reads
@@
For $k=1$, the $N\rightarrow 0$ result agrees with deGennes:1972zz . Moreover, the anomalous dimension of the symmetric traceless bilinear operator $\phi_{\text{ST}}^{2}$ is
$\gamma_{\phi_{\text{ST}}^{2}}=\frac{2}{N+8}\epsilon+O(\epsilon^{2})\,,$ (180)
@@
$D_{\text{F}}=d-\Delta_{\phi_{\text{ST}}^{2}}=2-\frac{2}{N+8}\epsilon+O(\epsilon^{2})\,,$ (181)
where the spacetime dimension is $d=4k-\epsilon$ and the fractal dimension $D_{\text{F}}$ is $k$-independent to order $\epsilon$. In particular, the limits $N\rightarrow 0$ and $N\rightarrow-2$ yield
@@
Let us further discuss the $n>2$ counterpart of the special $N$ limits. The $n>2$ theories are expected to be the multicritical generalizations of the critical O($N$) model at $d=4-\epsilon$. The tensor structures (173) are independent of $n$, so the $N\rightarrow 0$ limit is also singular in the case of $n>2$. The anomalous dimensions of $\phi$ read
$\gamma_{\phi}\big{|}_{N\rightarrow 0}$ $=\frac{(-1)^{k+1}(n-1)^{3}((n-1)!)^{2}\left(1+\frac{k}{n-1}\right)_{k}}{4k^{2}\,_{3}F_{2}\left(1-n,\frac{1}{2}-\frac{n}{2},-\frac{n}{2};1,\frac{1}{2}-n;1\right)^{2}\left(1+\frac{nk}{n-1}\right)_{k-1}(n)_{n}^{2}}\epsilon^{2}+O(\epsilon^{3})\,.$ (183)
The scaling dimensions of bilinear operators in the singlet and symmetric traceless representations are degenerate:
$\Delta_{{}_{\mathcal{J}_{\text{S}}}}\big{|}_{N\rightarrow 0}=\;$ $\Delta_{{}_{\mathcal{J}_{\text{ST}}}}\big{|}_{N\rightarrow 0}$
=\;= $d-2k+2m+\ell+\frac{((n-1)!)^{2}(n-1)^{2}\left(1+\frac{k}{n-1}\right)_{k}}{2(n)_{n}^{2}\,{}_{3}F_{2}\left(1-n,\frac{1}{2}-\frac{n}{2},-\frac{n}{2};1,\frac{1}{2}-n;1\right)^{2}}$
×$\times\left(\frac{(-1)^{k+1}(n-1)}{k^{2}\left(1+\frac{nk}{n-1}\right)_{k-1}}-\frac{(k-1)!(2n-1)\left(1+\frac{k}{n-1}\right)_{k}}{(k)_{k-m}(-m)_{m}\left(\ell+m-\frac{(n-2)k}{n-1}\right)_{2k}}\right)+O(\epsilon^{3})\,.$ (184)
Besides the singular tensor structure, we can identify more special $N$ limits by requiring that the bilinear operators are degenerate. There are three possibilities:
  • $\Delta_{{}_{\mathcal{J}_{\text{S}}}}|_{N\rightarrow 0}=\Delta_{{}_{\mathcal{J}_{\text{ST}}}}|_{N\rightarrow 0}$
These correspond to the case of singular tensor structure above. $\Delta_{{}_{\mathcal{J}_{\text{S}}}}|_{N\rightarrow-2}=\Delta_{{}_{\mathcal{J}_{\text{A}}}}|_{N\rightarrow-2}=d-2k+2m+\ell+O(\epsilon^{3})$ Here A denotes the antisymmetric representation. These degeneracies occur due to the vanishing of anomalous dimensions. $\Delta_{{}_{\mathcal{J}_{\text{all three reps}}}}|_{N\rightarrow-4,-6,\ldots,-2n+2}=d-2k+2m+\ell+O(\epsilon^{3})\qquad(n>2)$ These degeneracies are similar to those in the $N\rightarrow-2$ case, except that here all three types of bilinear operators are degenerate. We emphasize that the spin $\ell$ is arbitrary, so all the degeneracies above involve the entire Regge trajectories. The latter two scenarios also lead to $\gamma_{\phi}|_{N\rightarrow-2,-4,\ldots,-2n+2}=O(\epsilon^{3})\,.$ (185) We also comment on the Chew-Frautschi plots. In the $k=1$ case, the positive vertical-axis intercepts associated with $\Delta=d/2$ are ±$\ell^{*}_{\pm}=\frac{n-2}{n-1}+\left(\frac{1}{2}\pm\frac{n\sqrt{r}}{2^{n-3}(n)_{n}\,{}_{3}F_{2}\left(\frac{1}{2}-\frac{n}{2},-\frac{n}{2},-n-\frac{N}{2}+1;1,\frac{1}{2}-n;1\right)}\right)\epsilon+O(\epsilon^{2})\,,$ (186) where the ratio of OPE coefficients in the O($N$)-symmetric $\phi^{2n}$ theory is $r=\begin{cases}4^{n-1}(2n-1)(n-1)!\left(\frac{N}{2}+1\right)_{n-1}&\quad\!\!\!\text{singlet}\\ 2^{2n-3}(4n+N-2)(n-1)!\left(\frac{N}{2}+2\right)_{n-2}&\quad\!\!\!\text{symmetric traceless}\\ 4^{n-1}(n-1)!\left(\frac{N}{2}+1\right)_{n-1}&\quad\!\!\!\text{anti-symmetric.}\end{cases}$ (187) From this perspective, the limit $N\rightarrow-2$ is special in the sense that the two intercepts $\ell^{*}_{\pm}$ become identical for the singlet and antisymmetric representations respectively, and the limits $N\rightarrow-4,\ldots,-2n+2$ are special because $\ell^{*}_{+}=\ell^{*}_{-}$ for all three representations. For $k>1$, there are three qualitative differences. First, the Chew-Frautschi plots contain more curves. Second, there exist trajectories associated with higher $m$. Third, for each $m$, the equation $\Delta=d/2$ has more solutions, indicating that more intersections of the trajectory with the vertical axis can be found. These features are similar to those in the generalized Potts model, which are discussed at the end of subsection 3.1.